OpenStudy (anonymous):

simplify the radical, radical(4t^2+4t^4+1)

OpenStudy (dumbcow):

Factor :)

OpenStudy (anonymous):

\[ 4t^2+4t^4+1=4t^4+2t^2+2t^2+1=2t^2(2t^2+1)+1(2t^2+1)=(2t^2+1)^2 \]

OpenStudy (anonymous):

^ he is right. Another solution would be to represent t^2 as a It will simplify the equation to 4a^2 + 4a + 1 the factor would be (2a+1)^2 so substitute the real value of a which is t^2 it will be (2t2+1)2

OpenStudy (anonymous):

you guys are genius

OpenStudy (anonymous):

What was your final answer?

OpenStudy (anonymous):

well the question was to find the tangential unit vector, r(t)=<t^2,2/3 t^3, t>

OpenStudy (anonymous):

Oh I see.

OpenStudy (anonymous):

Using it for a line integral, or just something else?

OpenStudy (anonymous):

forgot about factoring lol. uhm, curvature. its a quiz problem says to find t(t) and n(t)